# A unified relative entropy framework for macroscopic limits of Vlasov–Fokker–Planck equations

### v1

@@ 1.2.2. High-field limit
$\|f^{\varepsilon}-M_{\rho,u}\|_{L^{\infty}(0,T;L^{1})}^{2}+\|\rho_{f^{\varepsilon}}-\rho\|_{L^{\infty}(0,T;L^{1})}^{2}+\|\rho_{f^{\varepsilon}}u_{f^{\varepsilon}}+\rho\nabla(-\Delta)^{-\alpha}\rho\|_{L^{\infty}(0,T;L^{1})}^{2}$
@@
If the initial data is mildly well-prepared, meaning that
@@
Remark 1.6.
In the mildly well-prepared case for the high-field limit, the initial relative entropy
$\mathscr{H}[f^{\varepsilon}|M_{\rho,-\nabla(-\Delta)^{-\alpha}\rho}](0)$
is not required to vanish as $\varepsilon\to 0$. Our framework remains valid even if it diverges, provided its growth is bounded by
@@
This contrasts with the diffusive regime, where the dissipation scales like $\frac{1}{\varepsilon^{2}}$ and thus allows initial defects as large as $\frac{1}{\varepsilon^{2-}}$. By comparison, the high-field scaling only involves factors of $\frac{1}{\varepsilon}$, leading to a stricter threshold. Moreover, the potential energy discrepancy must converge,
$\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)\to 0\quad\mbox{as }\varepsilon\to 0,$
since any growth in this term would be amplified by the prefactor $\frac{1}{\varepsilon}$ and prevent quantitative control.
We also provide an example of initial data $f^{\varepsilon}_{0}$ that satisfies the hypotheses of Theorem 1.2 and achieves the growth in (1.10). The construction parallels Remark 1.2. Consider
@@
$\varepsilon\mathscr{K}[f^{\varepsilon}_{0}]+\mathscr{P}[\rho_{f^{\varepsilon}_{0}}]=\mathscr{P}[\rho_{0}]+\frac{\varepsilon^{\delta}}{2}\int_{\mathbb{R}^{d}}\rho_{0}|v|^{2}\,dx<\infty$
@@
for sufficiently small $\varepsilon\in(0,1)$ and any $\delta\in(0,1)$, which yields (1.10).
In summary, the different admissible growth rates of $\mathscr{H}$ and $\mathscr{P}$ reflect the underlying scaling of the two limits: diffusive scaling permits larger initial entropy defects, while high-field scaling enforces a more restrictive preparation.
Remark 1.7.
Compared with the diffusive limit, where we were unable to obtain a quantified convergence result for the rescaled momentum $\frac{1}{\varepsilon}\rho^{\varepsilon}u^{\varepsilon}$ in the midly well-prepared case, the high-field scaling allows us to establish full quantitative control, including stability bounds for the momentum variable. This improvement reflects the fact that the suppression of kinetic transport in the high-field regime simplifies the structure of the limit equation, reducing it to the pure aggregation equation (1.9). The convergence and stability estimates thus highlight the effectiveness of combining the relative entropy method with BL estimates under both well-prepared and mildly well-prepared initial data.
@@
$\varepsilon\partial_{t}f^{\varepsilon}+\xi\cdot\nabla_{x}f^{\varepsilon}-\nabla(-\Delta)^{-\alpha}\rho_{f}^{\varepsilon}\cdot\nabla_{\xi}f^{\varepsilon}+\frac{1}{\varepsilon}\xi^{\perp}\cdot\nabla_{\xi}f^{\varepsilon}=\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon}).$ (1.11)
@@
If the initial data is mildly well-prepared in the sense that
${\rm d}_{\rm BL}(\rho_{f^{\varepsilon}_{0}},\rho_{0})\to 0\quad\mbox{and}\quad\varepsilon\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)\right)\to 0\quad\mbox{as }\varepsilon\to 0,$
then we have
$f^{\varepsilon}$ $\overset{*}{\rightharpoonup}M_{\rho,0}\quad\mbox{in }L^{2}(0,T;\mathcal{M}(\mathbb{R}^{2}\times\mathbb{R}^{2})),$
$\int_{\mathbb{R}^{2}}f^{\varepsilon}\,d\xi$ $\overset{*}{\rightharpoonup}\rho\quad\mbox{in }L^{\infty}(0,T;\mathcal{M}(\mathbb{R}^{2})).$
In this case, the following stability estimate holds:
$\sup_{0\leq t\leq T}{\rm d}^{2}_{\rm BL}(\rho_{f^{\varepsilon}}(t),\rho(t))+\int_{0}^{T}{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,0})\,dt$
$\quad\leq C{\rm d}^{2}_{\rm BL}(\rho_{f^{\varepsilon}_{0}},\rho_{0})+C\varepsilon\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)\right)+C\varepsilon^{2}.$
@@
In the mildly well-prepared case for the gSQG scaling, the convergence result still holds without requiring the initial modulated energies to vanish. In fact, one may allow
$\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)\sim\frac{1}{\varepsilon^{1-}}\quad\mbox{as }\varepsilon\to 0.$ (1.14)
This is less restrictive than in the high-field regime, where the potential part $\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)$ must converge to zero, but more restrictive than in the diffusive regime, which allows growth up to order $\frac{1}{\varepsilon^{2-}}$.
To provide a concrete instance, fix $\delta\in(0,1)$ and $v\in L^{2}(\rho_{0}\,dx)$, and set
$f^{\varepsilon}_{0}(x,\xi)=\frac{\rho_{0}(x)}{(2\pi)^{\frac{d}{2}}}\exp\left(-\frac{|\xi-u_{\varepsilon}(x)|^{2}}{2}\right),\qquad u_{\varepsilon}(x)=\frac{v(x)}{\varepsilon^{\frac{1-\delta}{2}}},$
with $\rho_{0}\in\mathcal{P}\cap\dot{H}^{-\alpha}(\mathbb{R}^{2})$. Then, by using almost the same arguments as in Remarks 1.2 and 1.6, one verifies that $f^{\varepsilon}_{0}$ satisfies the scaled initial bound in Theorem 1.3 and realizes the growth in (1.14).
Remark 1.10.
A key structural difference from the diffusive limit is the form of the kinetic corrector. In the gSQG scaling the corrector is $\nabla^{\perp}\rho$, which is divergence-free; hence, when testing the momentum balance against BL vector fields, the corrector cancels out upon taking divergence. This cancellation is precisely what allows us to close the quantitative estimate (1.13). By contrast, in the diffusive regime the corrector is $\nabla\rho$, which is curl-free and produces $\nabla\cdot(\nabla\rho)=\Delta\rho$ when paired with divergence, thereby producing the diffusion and preventing an analogous BL control of $\frac{1}{\varepsilon}\rho_{f^{\varepsilon}}u_{f^{\varepsilon}}$ within our framework. Accordingly, for the diffusive limit we only obtain quantitative control for the density ((in $L^{1}\cap\dot{H}^{-\alpha}$)), while an estimate for the rescaled momentum analogous to the gSQG case does not seem immediate within our approach and would require further investigation.
@@ 1.3. Novelty and contributions
The present work provides the first unified analysis of three representative asymptotic regimes for the Vlasov–Fokker–Planck equation with Riesz interactions: the diffusive limit, the high-field limit, and the gSQG limit. Our approach is based on a common relative entropy framework, supplemented with modulated interaction energies and bounded Lipschitz stability estimates. This combination yields not only qualitative convergence but also explicit quantitative error bounds, and allows us to treat both well-prepared and mildly well-prepared initial data, even without a uniform bound on the total energy.
@@ 1.3.1. Diffusive limit
where ${\rm d}_{2}$ denotes the second-order Wasserstein distance, so that the convergence rate is saturated by the $\varepsilon$-term regardless of the accuracy of the initial data, our analysis shows that for suitably prepared initial conditions the error can be reduced to $O(\varepsilon^{2})$. We also would like to remark that the $\varepsilon$ bound also appeared even in the perturbative framework [61]. From the perspective of time scaling, this improved rate can be regarded as essentially optimal.
@@ 1.3.2. High-field limit
Our contribution is to introduce a genuinely quantitative framework for the high-field limit with diffusion. By employing relative entropy estimates directly at the kinetic level, we exploit the dissipative structure of the Fokker–Planck operator and derive explicit convergence rates under well-prepared initial data. This represents, to our knowledge, the first quantitative justification of the high-field regime for Vlasov–Fokker–Planck-type models.
@@ 1.4. Organization of paper
The rest of this paper is organized as follows. In Section 2, we develop the analytic framework underlying our asymptotic analysis. This includes structural decompositions of the relative entropy and uniform energy–dissipation identities. Section 3 establishes quantitative stability tools, combining relative entropy, bounded Lipschitz distances, and modulated interaction energies. In Section 4, these methods are applied to derive the hydrodynamic limits in the diffusive, high-field, and gSQG regimes. Finally, Appendices provide a formal Hilbert expansion for the gSQG scaling and verify the propagation of regularity assumptions for the limiting equations.
@@ 2. Preliminaries
In this section, we present the analytic framework that underpins our asymptotic analysis. The discussion is organized around two complementary ingredients. We first recall structural decompositions and functional inequalities that allow entropy methods to connect microscopic information with macroscopic quantities. We then establish energy–dissipation balances that provide uniform control of kinetic and interaction energies throughout the evolution. Taken together, these tools form the basis for the quantitative convergence estimates developed in the next section.
@@ 2.1. Relative entropy and fundamental properties
We begin with the concept of relative entropy with respect to local Maxwellians. This quantity plays a central role in our analysis since it naturally decomposes into microscopic deviations and macroscopic mismatches. In addition, several classical inequalities link the relative entropy to Fisher information, $L^{1}$-type distances, and coercivity estimates. These results will serve as the starting point for the stability and convergence estimates established in later sections.
@@
Besides the structural decomposition above, several classical inequalities for the relative entropy will be frequently used in our analysis. They connect entropy with Fisher information, provide coercivity in the density variable, and relate entropy to $L^{1}$-type distances between distributions. For completeness, we record them here.
@@
Under the same assumptions as Lemma 2.1, the followings hold:
@@ 2.2. Energy inequalities
Strightforward computations give
@@
3. Quantitative stability via relative entropy, BL distance, and modulated interaction energy
In this section, we develop quantitative stability tools that link the dissipation structure of the kinetic equation with weak topologies suited for convergence analysis. The energy estimates in Section 2 already provide robust control of kinetic fluctuations around their associated local Maxwellians, and the crucial point is that the dissipation terms generated in the relative entropy identity carry scaling prefactors of order $\frac{1}{\varepsilon}$ or $\frac{1}{\varepsilon^{2}}$. The key coercive term here is the Fisher information dissipation, which directly controls microscopic deviations from the local Maxwellian. This amplification enables us to propagate stability even under mildly well-prepared initial data, where only partial compatibility with the limit profile is assumed. To make this effective, we complement the relative entropy framework with stability tools in weak topology, starting with bounded Lipschitz estimates and then turning to modulated interaction energies. These will allow us to convert microscopic dissipation into quantitative convergence toward the macroscopic system.
3.1. Bounded Lipschitz distance: connection with entropy and dissipation
As a first step, we analyze the BL distance. While the relative entropy provides a natural measure of distance to local Maxwellians, many of our convergence statements will be formulated in the weak topology of probability measures. For this purpose, the BL distance is particularly useful, since it metrizes weak convergence on $\mathcal{P}(\mathbb{R}^{d})$ (and on $\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})$) without additional moment assumptions; on bounded domains it is equivalent to the first-order Wasserstein distance up to constants. We next provide estimates that relate the BL distance to modulated kinetic quantities.

### v2

@@ 1.2.2. High-field limit
$\|f^{\varepsilon}-M_{\rho,-\nabla(-\Delta)^{-\alpha}\rho}\|_{L^{\infty}(0,T;L^{1})}^{2}+\|\rho_{f^{\varepsilon}}-\rho\|_{L^{\infty}(0,T;L^{1})}^{2}+\|\rho_{f^{\varepsilon}}u_{f^{\varepsilon}}+\rho\nabla(-\Delta)^{-\alpha}\rho\|_{L^{\infty}(0,T;L^{1})}^{2}$
@@
If the initial data is mildly well-prepared in the sense that
@@
The notion of mildly well-prepared initial data can be interpreted as an intermediate regime between the classical well-prepared and the fully ill-prepared settings. In contrast to the standard well-prepared case, we do not require the relative entropy $\mathscr{H}[f^{\varepsilon}|M_{\rho,-\nabla(-\Delta)^{-\alpha}\rho}](0)$ to vanish as $\varepsilon\to 0$. Our framework permits substantially larger deviations of the kinetic initial state from equilibrium, and in fact accommodates situations where this quantity diverges, with growth up to order
@@
To illustrate this, consider
@@
$\varepsilon\mathscr{K}[f^{\varepsilon}_{0}]+\mathscr{P}[\rho_{f^{\varepsilon}_{0}}]=\mathscr{P}[\rho_{0}]+\frac{\varepsilon^{\delta}}{2}\int_{\mathbb{R}^{d}}\rho_{0}|v|^{2}\,dx+\frac{d}{2}\int_{\mathbb{R}^{d}}\rho_{0}\,dx<\infty$
@@
for sufficiently small $\varepsilon\in(0,1)$ and any $\delta\in(0,1)$, which yields (1.9). This shows that our convergence result remains valid well beyond the classical well-prepared regime, and encompasses a range of ill-prepared configurations with large initial modulated energies.
${\rm d}_{\rm BL}(\rho_{f^{\varepsilon}_{0}},\rho_{0})\to 0,$
which ensures that the macroscopic density is correctly aligned with the limiting initial data. This condition is essential for quantitative convergence estimates: without such consistency at the level of densities, one cannot close the stability inequalities. If this condition is removed, only qualitative convergence can be expected, typically via compactness arguments (see [54] for example).
This viewpoint is also consistent with the formal asymptotic expansion presented in Appendix B.2. In the high-field regime, the expansion suggests that both the kinetic solution and the macroscopic density converge at first order, namely
$f^{\varepsilon}-M_{\rho,-\nabla(-\Delta)^{-\alpha}\rho}=O(\varepsilon),\quad\rho_{f^{\varepsilon}}-\rho=O(\varepsilon),$
with no additional cancellation in the density variable. In this sense, the BL-based stability estimate in Theorem 1.2 is consistent with the formally optimal scaling in a weak topology. More precisely, whenever the right-hand side of the BL-based stability estimate is of order $O(\varepsilon^{2})$, the estimate yields the natural first-order convergence rate at the level of weak distances.
@@
$\varepsilon\partial_{t}f^{\varepsilon}+\xi\cdot\nabla f^{\varepsilon}-\nabla(-\Delta)^{-\alpha}\rho_{f^{\varepsilon}}\cdot\nabla_{\xi}f^{\varepsilon}+\frac{1}{\varepsilon}\xi^{\perp}\cdot\nabla_{\xi}f^{\varepsilon}=\nabla_{\xi}\cdot(\nabla_{\xi}f^{\varepsilon}+\xi f^{\varepsilon}).$ (1.10)
@@
If we further assume that $\alpha\in[\frac{1}{2},1)$ and $\rho\in L^{\infty}(0,T;H^{\gamma+1}(\mathbb{R}^{2}))$ for some $\gamma\geq 4$, then the density error also satisfies the following quantitative estimate in the negative Sobolev space $H^{-\gamma}(\mathbb{R}^{2})$:
$\|(\rho_{f^{\varepsilon}}-\rho)(t)\|_{H^{-\gamma}}\leq C\left(\|\rho_{f_{0}^{\varepsilon}}-\rho_{0}\|_{H^{-\gamma}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho_{f^{\varepsilon}}|\rho](0)+\varepsilon\right).$ (1.13)
@@
Formal computations in Appendices A and B show that
$f^{\varepsilon}=M_{\rho,0}+\varepsilon f_{1}+O(\varepsilon^{2}),$
with
$f_{1}=-\left((\nabla\rho+\rho\nabla(-\Delta)^{-\alpha}\rho)^{\perp}\cdot\xi\right)M_{1,0}.$
Since $f_{1}$ is odd in $\xi$, we have
$\int_{\mathbb{R}^{2}}f_{1}\,d\xi=0.$
Thus, at the formal level, this suggests
$f^{\varepsilon}-M_{\rho,0}=O(\varepsilon),\quad\rho_{f^{\varepsilon}}-\rho=O(\varepsilon^{2}).$
In particular, as in the diffusive regime, the density variable gains one additional order due to the oddness of the first-order corrector in the velocity variable. We emphasize, however, that Theorem 1.3 only yields an $O(\varepsilon)$-type quantitative stability estimate, and does not address this formally sharper $O(\varepsilon^{2})$ density rate.
On the other hand, the first-order correction can still be absorbed into a shifted Maxwellian. Indeed, introducing the effective velocity
${\rm u}_{\varepsilon}:=-\varepsilon\left(\nabla^{\perp}\log\rho+\nabla^{\perp}(-\Delta)^{-\alpha}\rho\right),$
one may rewrite the first-order term as
$\varepsilon f_{1}=\xi\cdot{\rm u}_{\varepsilon}\,M_{\rho,0}.$
Hence, using the same first-order Taylor expansion of $M_{\rho,{\rm u}_{\varepsilon}}$ around ${\rm u}=0$ as in the diffusive case, one formally obtains the improved approximation
$f^{\varepsilon}-M_{\rho,{\rm u}_{\varepsilon}}=O\bigl(\varepsilon^{2}(1+|\xi|^{2})M_{\rho,0}\bigr).$
Thus, while the reference Maxwellian $M_{\rho,0}$ approximates the kinetic distribution only at order $O(\varepsilon)$, the shifted Maxwellian $M_{\rho,{\rm u}_{\varepsilon}}$ captures the leading drift correction and yields a formally sharper second-order approximation at the kinetic level. A corresponding rigorous estimate for $f^{\varepsilon}-M_{\rho,{\rm u}_{\varepsilon}}$ in a weaker negative Sobolev topology is given in Remark 4.2.
A key structural feature of the gSQG scaling is that the first-order momentum flux contains the divergence-free component $\nabla^{\perp}\rho$. As a consequence, when the momentum balance is paired with gradients of test functions, this part disappears thanks to the identity $\nabla\cdot\nabla^{\perp}\equiv 0$. This cancellation underlies the weak estimate (1.12) for the rescaled momentum.
This mechanism should be contrasted with the other two regimes. In the diffusive limit, the parabolic structure of the limiting equation makes it possible to obtain sharper quantitative estimates for the density in weaker topologies, in a way consistent with the formally optimal $O(\varepsilon^{2})$ scaling. In the high-field regime, the BL framework is naturally adapted to the transport structure of the limit equation and leads to first-order weak convergence estimates consistent with the formal asymptotics.
By contrast, although the formal expansion in the gSQG regime also suggests the improved density scaling $\rho_{f^{\varepsilon}}-\rho=O(\varepsilon^{2})$ due to the oddness of the first-order corrector, this sharper rate is not directly captured by our rigorous argument. The reason is that the limiting equation is now a conservative active scalar equation with Hamiltonian structure, rather than a parabolic equation with regularizing effects. Accordingly, the natural quantitative object in this regime is different from those in the diffusive and high-field cases: rather than a refined density estimate of the same type, the relevant object is the rescaled momentum, for which the rotational cancellation encoded in $\nabla^{\perp}\rho$ yields the appropriate weak stability structure. In this sense, estimate (1.13) reflects the weak quantitative structure intrinsic to the gSQG scaling, although the formally suggested $O(\varepsilon^{2})$ density improvement remains beyond the scope of the present argument.
@@ 1.3. Novelty and contributions
The present work provides the first unified analysis of three representative asymptotic regimes for the Vlasov–Fokker–Planck equation with Riesz interactions: the diffusive limit, the high-field limit, and the gSQG limit. Our approach is based on a common relative entropy framework, supplemented with modulated interaction energies and bounded Lipschitz stability estimates. This combination yields not only qualitative convergence but also explicit quantitative error bounds. It also allows us to treat well-prepared initial data in all three regimes and, in the high-field regime, a class of mildly well-prepared initial data even without a uniform bound on the total energy.
@@ 1.3.1. Diffusive limit
where ${\rm d}_{2}$ denotes the second-order Wasserstein distance, so that the convergence rate is saturated by the $\varepsilon$-term regardless of the accuracy of the initial data, our analysis shows that, for suitably prepared initial data, the error can be reduced to $O(\varepsilon^{2})$ in low-regularity topologies. We also note that the perturbative analysis in [64] yields an $O(\varepsilon)$ convergence rate for the kinetic distribution $f^{\varepsilon}$, but does not address the improved convergence of the macroscopic density $\rho^{\varepsilon}$. In this sense, the $O(\varepsilon^{2})$ density estimate obtained here is consistent with the formally optimal diffusive scaling.
@@ 1.3.2. High-field limit
Our contribution is to introduce a genuinely quantitative framework for the high-field limit with diffusion. By employing relative entropy estimates directly at the kinetic level, we exploit the dissipative structure of the Fokker–Planck operator and derive explicit convergence rates under well-prepared initial data. This represents, to our knowledge, the first quantitative justification of the high-field regime for Vlasov–Fokker–Planck-type models. Moreover, in the weak-topology setting, our BL-based estimates are consistent with the first-order scaling predicted by the formal asymptotic expansion.
@@ 1.4. Organization of paper
The rest of this paper is organized as follows. In Section 2, we develop the analytic framework underlying our asymptotic analysis. This includes structural decompositions of the relative entropy and uniform energy-dissipation identities. Section 3 establishes quantitative stability tools, combining relative entropy, weak-topology estimates, and modulated interaction energies. In Section 4, these methods are applied to derive the hydrodynamic limits in the diffusive, high-field, and gSQG regimes. Finally, the appendices provide the formal derivation of the gSQG limit, a discussion of the formal optimal convergence rates in the three singular regimes, and the verification of the regularity assumptions imposed on the limiting equations.
@@ 2. Preliminaries
In this section, we present the analytic framework that underpins our asymptotic analysis. The discussion is organized around two complementary ingredients. We first recall the basic structural properties of relative entropy, including its variational characterization, decomposition formulas, and functional inequalities, which allow entropy methods to connect microscopic information with macroscopic quantities. We then establish energy-dissipation balances that provide uniform control of kinetic and interaction energies throughout the evolution. Taken together, these tools form the basis for the quantitative convergence estimates developed in the next section.
@@ 2.1. Relative entropy and fundamental properties
We begin with the concept of relative entropy with respect to local Maxwellians. This quantity plays a central role in our analysis: on the one hand, it admits a natural decomposition into microscopic deviations and macroscopic mismatches; on the other hand, it enjoys several useful variational and functional-analytic properties. In particular, classical inequalities relate relative entropy to Fisher information, $L^{1}$-type distances, and coercivity estimates.
A first important aspect is its variational interpretation. For prescribed $(\rho,u)$, the local Maxwellian $M_{\rho,u}$ is the natural equilibrium state associated with the mass and momentum constraints, and in fact arises as the unique minimizer of a constrained free energy functional. We recall this characterization here, since it clarifies the distinguished role of Maxwellians in the relative entropy method.
More specifically, $M_{\rho,u}$ is the unique minimizer of the variational problem:
This characterization can be justified in several ways:
gives the identity
$\log g(\xi)+1+\frac{1}{2}|\xi|^{2}+\lambda_{0}+\lambda\cdot\xi=0,$
which implies
Completing the square in the exponent,
so that
$g(\xi)=C_{0}e^{\frac{1}{2}|\lambda|^{2}}e^{-\frac{1}{2}|\xi+\lambda|^{2}}.$
$C_{0}e^{\frac{1}{2}|\lambda|^{2}}=\frac{\rho}{(2\pi)^{\frac{d}{2}}},$
$g(\xi)=M_{\rho,u}(\xi),$
which shows that the Maxwellian is the unique minimizer of (2.1).
Beyond this variational characterization, relative entropy also admits a useful structural decomposition into a microscopic part and a macroscopic part. This decomposition will be one of the main tools in our later stability estimates, since it allows us to separate the deviation from the local Maxwellian with the same moments as $f$ from the mismatch between those moments and a prescribed target state. We now make this structure precise.
@@
In addition to the decomposition above, several classical inequalities for the relative entropy will be frequently used in our analysis. They connect entropy with Fisher information, provide coercivity in the density variable, and relate entropy to $L^{1}$-type distances between distributions. For completeness, we record them here.
@@
Under the same assumptions as in Lemma 2.1, the following properties hold:
@@ 2.2. Energy inequalities
Straightforward computations give
@@
3. Quantitative stability via relative entropy, weak-topology estimates, and modulated interaction energy
In this section, we develop quantitative stability tools that link the dissipation structure of the kinetic equation with weak topologies suited for convergence analysis. The energy estimates in Section 2 already provide robust control of kinetic fluctuations around their associated local Maxwellians, and the crucial point is that the dissipation terms generated in the relative entropy identity carry scaling prefactors of order $\frac{1}{\varepsilon}$ or $\frac{1}{\varepsilon^{2}}$. The key coercive term here is the Fisher information dissipation, which directly controls microscopic deviations from the local equilibrium. Combined with weak-topology stability tools, this mechanism allows us to convert microscopic dissipation into quantitative convergence estimates for the macroscopic system, both in strong topologies under well-prepared data and, in suitable regimes, in weaker topologies adapted to low-regularity or mildly prepared settings.
3.1. Weak-topology estimates: BL distance and negative Sobolev norms
In this subsection, we derive weak-topology estimates that connect the entropy-dissipation structure of the kinetic equation with quantitative convergence tools. While the relative entropy provides a natural measure of distance to local Maxwellians, several of our convergence statements are formulated in weak topologies on probability measures. As a first step, we therefore analyze the bounded Lipschitz distance, which metrizes weak convergence on $\mathcal{P}(\mathbb{R}^{d})$ and $\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{d})$ without additional moment assumptions; on bounded domains, it is equivalent to the first-order Wasserstein distance up to constants. We then complement the BL framework with negative Sobolev estimates, which will be useful when converting microscopic dissipation into macroscopic stability estimates in weak spatial topologies.